# Asymptotic error distribution for stochastic Runge–Kutta methods of strong order one

### v1

@@ 4.2. Properties of limit equation
Next, we show that the SRK method (3.9) is of weak order $2$ provided $\eta_{1}=0$.
@@
Let Assumptions 2 hold. Assume that $f\in\mathbf{C}^{6}(\mathbb{R}^{d})$ and $g\in\mathbf{C}^{7}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}f,\,\mathcal{D}^{7}g\in\mathbf{F}$. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order $2$, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
@@
with $P_{0}=Y_{0}$. In [21], for the scalar noise case, it has been shown that $\{P_{n}\}_{n=0}^{N}$ converges to $Y$ if conditions [21, (1)-(17)] holds. Note that [21, (1)-(17)] is equivalent to $\eta_{1}=0$, $\alpha^{\top}e=\beta^{\top}e=1$, and $\beta^{\top}(Be)=\frac{1}{2}$. Thus, under the assumptions of the theorem,
@@
In order to prove $\{Y^{SRK}_{n}\}_{n=0}^{N}$ is of weak order $2$, it suffices to prove that $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ for some $\iota_{1}\geq 1$ in view of (4.23). Thus the proof is complete once we show that there is some $K(y)\in\mathbf{F}$ such that for $k=1,\dots,5$,
@@
Note that among all terms in the final expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ and $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, the expectations of those terms with strong order $0.5$, $1.5$ and $2.5$ will vanish due to symmetry of distributions of $\Delta\widehat{W}$ and $\Delta W$. Thus, (4.27) holds if the expectations of those terms of strong $1$ and $2$ in the expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ is $3$th order approximation for corresponding terms in the expansions of $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, which is further implied by
@@
Let the assumptions of Lemma 4.7 hold. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order $2$, and the limit distribution of $N(Y^{SRK}_{N}-Y(T))$ as $N\to\infty$ takes the following unified form
@@ 5. Asymptotic error distribution for additive noise
where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is Lipschitz continuous, $X_{0}$ is non-random, $\mathbf{W}=(W_{1}(t),\ldots,W_{m}(t))_{t\in[0,T]}$ is an $m$-dimensional standard Brownian motion, and $\sigma=(\sigma_{1},\ldots,\sigma_{m})\in\mathbb{R}^{d\times m}$ is a constant matrix.
@@
    The SRK method is of strong order $1$, $i.e.,$ for any $p\geq 1$,
@@
    If additionally assume that $\bar{\alpha}^{\top}(\bar{A}e)=\bar{\alpha}^{\top}\bar{b}^{2}=\bar{\alpha}^{\top}\bar{b}=\frac{1}{2}$, then then the SRK method (3.9) is of weak order $2$, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
@@
Following the argument for proving $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ in Lemma 4.7 (or using [17, Page 100,Theorem 2.1]), the third conclusion holds provided that for $k=1,\ldots,5$,
@@
Let Assumption 3 hold. Then $\{\widetilde{X}_{n}\}_{n=0}^{N}$ is of strong order $1.5$ and $1$ when approximating $\{X^{SRK}_{n}\}_{n=0}^{N}$ and $\{X(t)\}_{t\in[0,T]}$, respectively. More precisely, we have that for any $p\geq 1$,
@@
Further, it is not hard to verify that conditions of Theorem 2.1 are satisfied with $p_{1}=3$ and $p_{2}=2$. Thus, $\sup_{n=0,1,\ldots,N}\|\widetilde{X}_{n}-X^{SRK}_{n}\|_{\mathbf{L}^{p}(\Omega)}\leq Kh^{1.5}$. This, combined with Lemma 5.2(2) and the Minkowski inequality, completes the proof. ∎
@@
in $\mathbf{L}^{p}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$. Similar to the argument for $\|I_{3}^{N}(t)-I_{3}(t)\|_{\mathbf{L}^{p}(\Omega)}\to 0$ in the proof of Lemma 4.3, applying [5, Proposition 4.2] yields that $J_{2}^{N}(t)$ converges to
@@
[…]
@@
Let Assumption 3 hold. If $\eta_{2}=0$, then the SRK method (5.2) is of weak order $2$, and the limit distribution of $N(X^{SRK}_{N}-X(T))$ as $N\to\infty$ takes the following unified form
@@
As is mentioned in Remark 4.9, Theorems 5.5 and 5.6 indicate that $\eta_{2}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method (5.2). In addition, we obtain the following inferences: (1) For the SRK methods (5.2) of strong order $1$, the smaller $\eta_{2}$ is, the smaller the mean-square error is after a long time. (2) Among the SRK methods (5.2) of strong order $1$, those of weak order $2$, i.e., $\eta_{2}=0$, have the smallest mean-square errors after a long time.
@@ 6. Numerical experiments
In this section, we perform numerical experiments to verify our theoretical analysis. For convenience, in this section, we consider SDEs with additive noise. We verify the convergence in distribution of $N(X^{SRK}_{N}-X(T))$ in Example 6.1, and verify the inferences of Remark 5.8 in Example 6.2 by presenting the evolution of the mean-square error of numerical methods w.r.t time.
@@
with $V_{0}=0$. The expectations of $\varphi(N(X_{N}-X(T)))$ and $\varphi(V(T))$ are realized by using the Monte–Carlo method with $150000$ sample paths. Table 1 displays the error $\big{|}\mathbf{E}\varphi(N(X_{N}-X(T)))-\mathbf{E}\varphi(V(T))\big{|}$ for different step-sizes and test functions $\varphi$. We observe that this error tends to 00 as $h$ decreases. This verifies the conclusion of Theorem 5.5.
Table 1. Error between $\mathbf{E}\varphi(N(X_{N}-X(T)))$ and $\mathbf{E}\varphi(V(T))$ for different step-sizes and $\varphi$.
@@
In this experiment, we consider four numerical methods for approximating (6.3): the trapezoid method, the $\theta$-method with $\theta=\frac{1}{2}$, $\frac{\sqrt{2}}{2}$ and $\theta=1$. The $\theta$-method is a $1$-stage Runge–Kutta method taking of form (5.2), whose Butcher tableau is $\theta$ $\theta$ $1$ In addition, the $\frac{1}{2}$-method and $1$-method are indeed the midpoint method and the implicit Euler method, respectively. It is not difficult to compute $\eta_{2}$ for these four methods.
@@
First, we test the mean-square convergence orders for the above four methods. We set $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=1$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with $5000$ sample paths. It is observed in Figure 1 and Table 3 that all these four methods have first order mean-square convergence. The mean-square error between the numerical solution of the trapezoid method and $X(T)$ is smaller than that between the numerical solution of the midpoint method and $X(T)$, although they are quite close.
Figure 1. Mean square errors for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$.
@@
Figure 2. Weak errors for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$.
@@
Figure 3. Evolution of mean-square errors of the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method with $h=0.01$.
@@
Second, we test the weak convergence orders of these methods. We take $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=2^{-2}$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with $200000$ sample paths. Figure 2 indicates that the weak order of the trapezoid method is $2$ and the weak order of the other three methods is $1$.

### v2

@@ 4.2. Properties of limit equation
Next, we show that the SRK method (3.9) is of weak order 222 provided $\eta_{1}=0$.
@@
Let Assumptions 2 hold. Assume that $f\in\mathbf{C}^{6}(\mathbb{R}^{d})$ and $g\in\mathbf{C}^{7}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}f,\,\mathcal{D}^{7}g\in\mathbf{F}$. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order 222, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
@@
with $P_{0}=Y_{0}$. In [20], for the scalar noise case, it has been shown that $\{P_{n}\}_{n=0}^{N}$ converges to $Y$ if conditions [20, (1)-(17)] holds. Note that [20, (1)-(17)] is equivalent to $\eta_{1}=0$, $\alpha^{\top}e=\beta^{\top}e=1$, and $\beta^{\top}(Be)=\frac{1}{2}$. Thus, under the assumptions of the theorem,
@@
In order to show that $\{Y^{SRK}_{n}\}_{n=0}^{N}$ is of weak order 222, it suffices to prove that $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ for some $\iota_{1}\geq 1$ in view of (4.23). Thus the proof is complete once we show that there is some $K(y)\in\mathbf{F}$ such that for $k=1,\dots,5$,
@@
Note that among all terms in the final expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ and $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, the expectations of those terms with strong order 0.50.50.5, 1.51.51.5 and 2.52.52.5 will vanish due to symmetry of distributions of $\Delta\widehat{W}$ and $\Delta W$. Thus, (4.27) holds if the expectations of those terms of strong 111 and 222 in the expansions of $\prod_{j=1}^{k}\bar{\Delta}^{i_{j}}$ is 333th order approximation for corresponding terms in the expansions of $\prod_{j=1}^{k}\bar{\bar{\Delta}}^{i_{j}}$, which is further implied by
@@
Let the assumptions of Lemma 4.7 hold. If $\eta_{1}=0$, then the SRK method (3.9) is of weak order 222, and the limit distribution of $N(Y^{SRK}_{N}-Y(T))$ as $N\to\infty$ takes the following unified form
@@ 5. Asymptotic error distribution for additive noise
where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is Lipschitz continuous, $X_{0}$ is non-random, and $\sigma=(\sigma_{1},\ldots,\sigma_{m})\in\mathbb{R}^{d\times m}$ is a constant matrix.
@@
    The SRK method is of strong order 111, $i.e.,$ for any $p\geq 1$,
@@
    If additionally assume that $\bar{\alpha}^{\top}(\bar{A}e)=\bar{\alpha}^{\top}\bar{b}^{2}=\bar{\alpha}^{\top}\bar{b}=\frac{1}{2}$, then the SRK method (3.9) is of weak order 222, i.e., for any $\varphi\in\mathbf{C}^{6}(\mathbb{R}^{d})$ with $\mathcal{D}^{6}\varphi\in\mathbf{F}$,
@@
Following the argument for proving $|\mathbf{E}\varphi(Y^{SRK}_{t,y}(t+h))-\mathbf{E}\varphi(P_{t,y}(t+h))|\leq K(1+|y|^{\iota_{1}})h^{3}$ in Lemma 4.7 (or using [16, Page 100,Theorem 2.1]), the third conclusion holds provided that for $k=1,\ldots,5$,
@@
Let Assumption 3 hold. Then $\{\widetilde{X}_{n}\}_{n=0}^{N}$ is of strong order 1.51.51.5 and 111 when approximating $\{X^{SRK}_{n}\}_{n=0}^{N}$ and $\{X(t)\}_{t\in[0,T]}$, respectively. More precisely, we have that for any $p\geq 1$,
@@
Further, it is not hard to verify that conditions of Theorem 2.1 are satisfied with $p_{1}=3$ and $p_{2}=2$. Thus, $\sup\limits_{n=0,1,\ldots,N}\|\widetilde{X}_{n}-X^{SRK}_{n}\|_{\mathbf{L}^{p}(\Omega)}\leq Kh^{1.5}$. This, combined with Lemma 5.2(2) and the Minkowski inequality, completes the proof. ∎
@@
in $\mathbf{L}^{p}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$. Similar to the argument for $\|I_{3}^{N}(t)-I_{3}(t)\|_{\mathbf{L}^{p}(\Omega)}\to 0$ in the proof of Lemma 4.3, applying [6, Proposition 4.2] yields that $J_{2}^{N}(t)$ converges to
@@
Further, following the ideas for proving Lemma 4.4, one can show that $J_{1}^{N}\overset{stably}{\Longrightarrow}J_{1}$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$, where
@@
Let Assumption 3 hold. If $\eta_{2}=0$, then the SRK method (5.2) is of weak order 222, and the limit distribution of $N(X^{SRK}_{N}-X(T))$ as $N\to\infty$ takes the following unified form
@@
As is mentioned in Remark 4.9, Theorems 5.5 and 5.6 indicate that $\eta_{2}$ is the key parameter reflecting the growth rate of the mean-square error of the SRK method (5.2). In addition, we obtain the following inferences: (1) For the SRK methods (5.2) of strong order 111, the smaller $\eta_{2}$ is, the smaller the mean-square error is after a long time. (2) Among the SRK methods (5.2) of strong order 111, those of weak order 222, i.e., $\eta_{2}=0$, have the smallest mean-square errors after a long time.
@@ 6. Numerical experiments
In this section, we perform numerical experiments to verify our theoretical analysis. For convenience, in this section, we consider SDEs with additive noise. We verify the convergence in distribution of $N(X^{SRK}_{N}-X(T))$ in Example 6.1, and verify the inferences of Remark 5.8 in Example 6.2 by presenting the evolution of the mean-square error of numerical methods w.r.t. time.
@@
with $V_{0}=0$. The expectations of $\varphi(N(X_{N}-X(T)))$ and $\varphi(V(T))$ are realized by using the Monte–Carlo method with 150000150000150000 sample paths. Table 2 displays the error $\big{|}\mathbf{E}\varphi(N(X_{N}-X(T)))-\mathbf{E}\varphi(V(T))\big{|}$ for different step-sizes and test functions $\varphi$. We observe that this error tends to 0 as $h$ decreases. This verifies the conclusion of Theorem 5.5.
Table 2. Error between $\mathbf{E}\varphi(N(X_{N}-X(T)))$ and $\mathbf{E}\varphi(V(T))$ for different step-sizes and $\varphi$ for (6.1).
@@
In this experiment, we consider four numerical methods for approximating (6.3): the trapezoid method, the $\theta$-method with $\theta=\frac{1}{2}$, $\frac{\sqrt{2}}{2}$, 111. The $\theta$-method is a 111-stage Runge–Kutta method taking of form (5.2), whose Butcher tableau is $\theta$ $\theta$ 111 . In addition, the $\frac{1}{2}$-method and 111-method are indeed the midpoint method and the implicit Euler method, respectively. It is not difficult to compute $\eta_{2}$ for these four methods.
@@
First, we test the mean-square convergence orders for the above four methods. We set $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=1$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with 500050005000 sample paths. It is observed in Figure 1 and Table 4 that all these four methods have first order mean-square convergence. The mean-square error between the numerical solution of the trapezoid method and $X(T)$ is smaller than that between the numerical solution of the midpoint method and $X(T)$, although they are quite close.
Figure 1. Mean square errors for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$.
@@
Figure 2. $|\mathbf{E}\exp(-X_{N})-\mathbf{E}\exp(-X(T))|$ for the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) in the log-log scale for five different step-sizes $h=2^{-4},2^{-5},2^{-6},2^{-7},2^{-8}$
@@
Figure 3. Evolution of mean-square errors of the trapezoid method, midpoint method, $\frac{\sqrt{2}}{2}$-method and implicit Euler method applied to (6.3) with $h=0.01$.
@@
Second, we test the weak convergence orders of these methods. We take $x_{0}=1$, $\sigma_{1}=\sigma_{2}=1$ and $T=2^{-2}$. These methods are realized by applying the fixed point iteration with a tolerance error $10^{-10}$. The exact solution $X(T)$ are approximated by the trapezoid method with the small step-size $h=2^{-16}$. The expectation is approximately obtained based on the Monte–Carlo method with 200000200000200000 sample paths. Figure 2 indicates that the weak order of the trapezoid method is 222 and the weak order of the other three methods is 111.